OpenStudy (anonymous):

find the integral 0-1 of (xe)^-(x)^2 dx with derivatives and integrals of the exponential function, i am very lost please help, i know that i plugg in the 0 and 1 into u and dx but after that i am very confused and the answer is (e-1)/(2e)^2 help

OpenStudy (aum):

Is it \[\int\limits_{0}^{1}xe^{-x^2}dx\]

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

a nice u sub will do it in one step

OpenStudy (aum):

Let u = x^2 du = 2xdx xdx = du/2 when x = 0, u = 0 when x = 1, u = 1 \( \large \int\limits_{0}^{1}xe^{-x^2}dx = \frac 12\int\limits_{0}^{1}e^{-u}du\)

OpenStudy (anonymous):

or you could have made \(u=-x^2\) etc

OpenStudy (aum):

yes.

OpenStudy (anonymous):

i can see how the demoninator is 2e^2, but how is the numerator 1-e, kinda lost there

OpenStudy (anonymous):

do you take the derivative of e^-u?

OpenStudy (aum):

If the problem I wrote in my first reply is correct, the answer I am getting is (e - 1) / (2e).

OpenStudy (anonymous):

oh yes, that is it, mixed the two